A person stands directly in front of two speakers that are emitting the same pure tone. The person then moves to one side until no sound is heard. At that point, the person is 4.90 m from one of the speakers and 4.95 m from the other. What is the frequency (in Hz) of the tone being emitted? (Assume air temperature to be 20°C.)
1
Expert's answer
2019-04-30T09:47:10-0400

Where a person stands there is a minimum of sound interference. Minimum rule

"l_2-l_1=\frac{\lambda}{2}"

And from phase velocity of wave

"v=\lambda \nu"

we have frequency

"\nu=\frac{v}{2\cdot (l_2-l_1)}"

So,

"\nu=\frac{330}{2 \cdot 0.05}=3300 (Hz)"